OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve two-thirds minus h divided by two equals five-eighths? A. five-eighths plus five-eighths B. five-eighths plus two-thirds C. five-eighths minus five-eighths D. five-eighths minus two-thirds

OpenStudy (stamp):

\[\frac{2}{3}-\frac{h}{2}=\frac{5}{8}\]

OpenStudy (stamp):

Is this correct setup?